The length of a spring is 40cm and its constant is half a newton per centimeter. If we tie a weight of half a kilogram to the end of it and pull it up at a Acceleration of 4m/s², what will be the length of the spring?
1
Expert's answer
2020-12-21T11:01:10-0500

"l_0=0.4(m)"

"k=500(N\/m)"


"k\Delta x_0=mg\to \Delta x_0=mg/k=0.5\cdot9.8/500=0.0098(m)"


"k\Delta x=ma\to \Delta x=ma/k=0.004(m)"


The length of the spring


"l=l_0+\Delta x_0+\Delta x=0.4+0.0098+0.004=0.4138(m)=41.38(cm)" . Answer